OpenStudy (anonymous):

calculate the density of a rectangular solid which has a mass of 25.71g. it is 2.30cm long, 4.01 cm wide, and 1.82 cm high.

OpenStudy (jfraser):

density is the ratio of an object's mass to its volume. Find the volume of the box.

OpenStudy (anonymous):

\[ d = mass/volum\] change all values to standard metric system (i dono what the system is called exactly) i mean, g to kg, cm to m and so on... d= 0.02571/(0.023*0.0401*0.0182) d= 0.02571/16.78586*10^-6 \[d=1531.646 \space kg/m^3\]

OpenStudy (aravindg):

whats ur difficulty here?